Consider the two-variable model where you want to maximize 2x1 + x2. There are three constraints: x1 + x2 <= 2, 4x1 + x2 <= 4, and -x1 + x2 <= 1. Also, x1and x2 are constrained to be nonnegative. Which of the following is the optimal solution?

a. x1 = 2/3, x2 = 4/3
b. x1 = 0, x2 = 1
c. x1 = 1/2, x2 = 3/2
d. x1 = 2, x2 = 0
Answer #1

Option a is the correct answer

x1= 2/3 , x2 = 4/3

If you liked the answer please give an Up-vote, this will be quite encouraging for me, thank you!